Balance the following redox equation in
a. acidic solution
b. basic solution
MnO4- + SO2= SO42- + Mn2+
1
Expert's answer
2019-11-26T06:40:03-0500

Solution.

a.

"2MnO4^- + 5SO2 + 2H2O = 5SO4^{2-} + 2Mn^{2+} + 4H^+"

b.

"2MnO4^- + 5SO2 + 4OH^- = 5SO4^{2-} + 2Mn^{2+} + 2H2O"

Answer:

a.

"2MnO4^- + 5SO2 + 2H2O = 5SO4^{2-} + 2Mn^{2+} + 4H^+"

b.

"2MnO4^- + 5SO2 + 4OH^- = 5SO4^{2-} + 2Mn^{2+} + 2H2O"